4) (a) Ordinary conversation at 50cm corresponds to a sound level of 65db. What is the new sound level if 2 of John's friends join in? (b) If the temperature in John's church rises, how will that affect the pitch of organ pipes? Briefly explain.
Answer #1


a)
let I is the sound intensity due to one person.

beta = 10*log(I/Io)

65 = 10*log(I/10^-12)

6.5 = log(I/10^-12)

10^6.5 = I/10^-12

I = 10^(-12 + 6.5)

I = 3.162*10^-6 W/m^2

when 2 of John's friends join in, new sound intensity

I' = 3*I

= 3*3.162*10^-6

= 9.486*10^-6

new sound intensity level,

beta' = 10*log(I'/Io)

= 10*log(9.486*10^-6/10^-12)

= 69.8 dB <<<<<<<<<<<---------------------------------Answer

b) normal mode frequencies of open organ pipes,

fn = n*v/(2*L) (n = 1,2,3,4,5,6,7,....)

v ---> speed of sound
L --> length of the pipes

If the temperature increases speed of sound also increases. so, pitch(frequency) also increases.